OpenStudy (anonymous):

Help with math please! find the value of x in a triangle with length of 8, height of 6 & diagonal of 5x-4

OpenStudy (anonymous):

i assume it is a right triangle?

OpenStudy (anonymous):

sorry i meant rectangle

OpenStudy (anonymous):

if so you have a 6 - 8 - 10 right triangle, so \(5x-4=10\)

OpenStudy (anonymous):

ok same thing

OpenStudy (anonymous):

Rectangle***

OpenStudy (anonymous):



OpenStudy (anonymous):

we know the diagonal is 10 by pythagoras \[6^2+8^2=10^2\]

OpenStudy (anonymous):

how do you get 10?

OpenStudy (anonymous):

pythagoras

OpenStudy (anonymous):

how do you get x though

OpenStudy (anonymous):

the diagonal makes a right triangle

OpenStudy (anonymous):

solve \(5x-4=10\) in two steps

OpenStudy (anonymous):

add 4, divide by 5